# Exact Screening-Ranged Expansions for Many-Body Electrostatics

### v1

@@ Abstract
We present an exact many-body framework for electrostatic interactions among $N$ arbitrarily charged spheres in an electrolyte, modeled by the linearized Poisson–Boltzmann equation. Building on a spectral analysis of nonstandard Neumann–Poincaré-type operators introduced in a companion mathematical work [supplem_pre_math], we construct convergent screening-ranged series for the potential, interaction energy, and forces, where each term is associated with a well-defined Debye–Hückel screening order and can be obtained evaluating an analytical expression rather than numerically solving an infinitely dimensional linear system. This formulation unifies and extends classical and recent approaches, providing a rigorous basis for electrostatic interactions among heterogeneously charged particles (including Janus colloids) and yielding many-body generalizations of analytical closed-form results previously available only for two-body systems. The framework captures and clarifies complex effects such as asymmetric dielectric screening, opposite-charge repulsion, and like-charge attraction, which remain largely analytically elusive in existing treatments. Beyond its fundamental significance, the method leads to numerically efficient schemes, offering a versatile tool for modeling colloids and soft/biological matter in electrolytic solution.
@@
For a single charged dielectric sphere in an electrolyte (the classical Kirkwood sphere [Kirkwood1934]) the LPBE admits a closed-form solution, but already for two interacting spheres ($N=2$) with different dielectric constants mutual polarization makes the problem highly nontrivial [Fish, Derb2, BesleyACR, LCSB, our_jcp, McClurg, Yu3]. Already in the Poisson limit of zero ionic strength, coupled multipolar expansions subject to infinitely many boundary constraints on the coefficients severely hinder analytical progress [GWJXL, our_jpcb, LCSB, BSL2014, Yu2019, DGC_softmat]. As emphasized in Ref. [BSL2014], “An implicit series expansion is known for the system of two dielectric spheres. For more than two dielectric spheres, numerical treatment is required.” Since then, image-based multiple-scattering formalisms [Qin2019, GXFQ, QPF, Freed2014], Wigner-matrix-free approaches [Yu2019, Yu3, Yu2021], and hybrid analytical-numerical schemes [GJLX, GWJXL, DGC_softmat, LSB2018, LinQS, HasStammI, Hassan_Stamm_jctc, Lindgren_jcis, FBYJBH] have substantially extended what can be done for two spheres, but a general explicit many-body ($N>2$) treatment at finite ionic strength is still lacking. Ionic screening further increases the technical complexity [Fish, LLF, LinQS, DuanGan2025, Yu3], limits the applicability of image-charge constructions [DuanGan2025], and leaves key questions on like-charge attraction (LCA), opposite-charge repulsion (OCR), and asymmetric dielectric screening (ADS) largely unresolved [Yu2021, Derb2, LCSB].
@@
In this Letter, we address this gap within the LPBE/DH framework by developing an exact and explicit many-body formalism for $N$ arbitrarily charged dielectric spheres in an electrolyte. Using the spectral properties of a nonstandard Neumann–Poincaré-type operator introduced in the companion mathematical work [supplem_pre_math], we construct a convergent Neumann-type expansion for the multipole coefficients of the DH potentials and thereby obtain screening-ranged series for the potentials, interaction energies, and forces. Each term in these series is associated with a definite “screening order” and contains products of Debye–Hückel factors $e^{-\kappa R_{ij}}/R_{ij}$ along interaction paths, providing a systematic and controllable many-body extension of the familiar DLVO and unscreened Coulomb descriptions. The formalism applies to general fixed charge distributions, including heterogeneously charged and Janus-like particles, and yields explicit expressions that are straightforward to evaluate numerically.
@@
We illustrate the scope of our formalism by (i) deriving exact higher-order screened corrections for charged spheres, clarifying the conditions for LCA, OCR, and ADS, and (ii) constructing compact DLVO-like interaction expressions for Janus particles with controlled many-body and higher-multipole corrections. Then, we highlight its value through two representative applications. For monopolar charged spheres, we recover the standard DLVO interaction as the leading screening order and obtain exact higher-order many-body corrections that generalize two-sphere Fisher–Levin–Li, McClurg–Zukoski, and related energy and force expressions [Fish, LLF, McClurg, FinBar2016, our_jcp, Yu3, Derb2, Derb4, BesleyACR, FilStar_jpcb]. Even in the long-studied two-sphere problem, this framework yields new analytical insight: it provides explicit conditions for the (non)occurrence of LCA and OCR, clarifies the role of dielectric contrast and ionic strength in ADS, and recovers recent image-charge-based polarization-energy results as a particular instance of the double-screened contribution [DGC_softmat, DuanGan2025, RPSA_2024, Chan]. For Janus-type particles with bipatchy surface charge, we derive compact DLVO-like expressions for effective monopole-monopole, monopole-dipole, and dipole-dipole interactions [GraafBoon2012, GLCBB_2025], and show how higher screening orders systematically generate many-body and higher-multipole corrections. Taken together, these examples demonstrate that the proposed screening-ranged expansions provide a versatile analytical framework for electrostatic interactions in model colloids and coarse-grained biomolecular systems, directly amenable to efficient numerical implementation. Beyond the specific cases discussed here, the framework offers a controlled many-body expansion for screened electrostatics, organized in a compact, rapidly convergent series that can underpin effective interaction models in charged soft-matter systems, from patchy colloids to biomolecular assemblies; further examples, including particles with clouds of internal point charges, higher-order multipoles, or inhomogeneous free surface distributions, are developed in the joint papers [supplem_pre, supplem_pre_force].
@@
where $\vec{\tilde{\mathbb{G}}}$ collects all exterior multipole coefficients, $\vec{\tilde{\mathbb{S}}}$ gathers the source terms due to fixed charges, $\mathbb{I}$ is the identity and $\mathbb{K}$ is a compact operator that couples different particles and multipoles through screened propagators proportional to $e^{-\kappa R_{ij}}/R_{ij}$ ($R_{ij}$ is the magnitude of vector $\mathbf{R}_{ij}=\mathbf{x}_{j}-\mathbf{x}_{i}$ connecting sphere centers). A detailed construction of $\mathbb{K}$ and $\vec{\tilde{\mathbb{S}}}$ is given in [supplem_pre_math, supplem_pre] (see End Matter). The spectral analysis in [supplem_pre_math] shows that the spectral radius of $\mathbb{K}$ considered in the relevant Hilbert space of multipole coefficients is strictly less than $1$. This guarantees that the inverse of $(\mathbb{I}+\mathbb{K})$ admits a convergent Neumann-type operator series representation,
@@
The resulting screening-ranged many-body energy and force expansions (4)-(5) converge rapidly with $\ell$ [supplem_pre, supplem_pre_force] and are easy to implement numerically. Thus, the first few addends are typically sufficient to accurately approximate the total energy $\mathcal{E}$ and force $\mathbf{F}_{i}$, and avoid the need for conventional approaches that require solving laborious systems of equations for potential coefficients or surface polarization charges. As documented in the companion works [supplem_pre, supplem_pre_force], this rapid convergence holds for representative colloidal and biomolecular parameter sets. Since the most computationally demanding step in harmonic-expansion-based approaches is solving the linear system / inverting the corresponding matrix [DY_2006], the identification of easy-to-calculate first addends of (4)-(5) already constitutes a major computational simplification. In the companion papers [supplem_pre, supplem_pre_force], we present several examples and show that including up to the double-screened addends ($\ell=2$) already captures the key features of the full energy and force profiles, including LCA and OCR.
@@
Importantly, this framework allows identifying the specific conditions for the (non)occurrence of less obvious phenomena, such as OCR and LCA. Indeed, in the particular case of two spheres less polarizable than the surrounding medium ($\varepsilon_{i}\leq\varepsilon_{\text{sol}}$) and arbitrary $q_{i}$, it turns out that odd-$\ell$ addends $\mathcal{E}^{(\ell)}\bigr|_{N=2}(R)$, such as DLVO ($\ell=1$), have the usual Coulombic behavior w.r.t. $R$, showing opposite-charge attraction and like-charge repulsion, while even-$\ell$ addends with $\ell\geq 2$ are repulsive irrespective of the charge. This structure permits OCR (through the competition between attractive odd-$\ell$ and repulsive even-$\ell$ contributions) [^2] Indeed, when $q_{1}q_{2}<0$, competition between attractive odd-$\ell$ and repulsive even-$\ell$ addends can produce short-range OCR [supplem_pre]. This corresponds to the asymmetric dielectric screening effect observed in numerical modelling in [DY_2006, Yu2021]. but forbids LCA. However, this is not the case when particles are more polarizable than the medium ($\varepsilon_{i}>\varepsilon_{\text{sol}}$) and $\kappa>0$, where the ionic strength can play a decisive role on the occurrence of OCR and LCA. Despite the existing recent DH-level treatments on the topic (see e.g. [Yu2021] and references therein), to our knowledge, OCR in this regime has not been explicitly characterized before in the literature (see [supplem_pre] for technical details).
@@
Figure 1: a) Electrostatic potential and iso-potential contour lines obtained for a three-sphere system. $\varepsilon_{1-3}=2$, $\varepsilon_{\text{sol}}=80$, $q_{1}=3e$, $q_{2,3}=-2e$, $\kappa^{-1}=8.07$ Å. The arrows represent the forces acting on the spheres. Spheres 22 and $3$ are identical, but located at a different distance from sphere 1. The most significant correction introduced by the inclusion of higher order terms consists in an excess negative potential located inside sphere 3. This potential is more pronounced in the region facing sphere 1. This induces a difference in the particle polarization, reflected by the difference in surface charge distribution shown in panel b). c) Two spheres interaction energy w.r.t. $R=R_{12}$, $R\geq 36$ Å; $a_{1}+a_{2}=35$ Å, $a_{1}=10a_{2}$, $\varepsilon_{1}=2$, $\varepsilon_{2}=3$, $q_{1}=3e$, $q_{2}=-2e$, $\varepsilon_{\text{sol}}=80$, $\kappa^{-1}=40$ Å. Lines F, 1 and 2 depict the full $\mathcal{E}^{\text{Int}}$, $\mathcal{E}^{(1)}$ (i.e. DLVO) and $\mathcal{E}^{(1)}+\mathcal{E}^{(2)}$, respectively. Lines 2.0, 2.1 and 2.1’ correspond to one-term (i.e. MZ), two-term and simplified two-term (i.e. FLL) approximations of $\mathcal{E}^{(2)}$, respectively. (Embedded inset shows a close-up view). d) Example of $\sigma_{i}^{\text{f}}$ for a generalized Janus particle: free charges $q_{i,1}$ and $q_{i,2}$ are uniformly distributed over spherical caps of polar angles $0,\,\theta_{i,1}]$ and $\pi-\theta_{i,2},\,\pi]$, respectively, while the intermediate surface (i.e. $(\theta_{i,1},\,\pi-\theta_{i,2})$) has no fixed charge.
@@
In summary, we have introduced an exact many-body framework for electrostatics of dielectric spheres in electrolytic solution within the LPBE/DH description. By exploiting the spectral properties of a Neumann–Poincaré-type operator introduced in [supplem_pre_math], we obtain convergent screening-ranged expansions for potentials, energies, and forces, organized by products of DH screening factors and directly applicable to arbitrary fixed charge distributions. The formalism recovers familiar DLVO interactions at leading order and provides explicit higher-order corrections that quantify many-body and mutual-polarization effects, clarify the conditions for ADS and LCA/OCR, and connect to recent image-charge and effective-dipole results [Fish, LLF, FinBar2016, our_jcp, DGC_softmat, DuanGan2025, RPSA_2024, Chan, Qin2019, GXFQ, QPF, Freed2014]. Our treatment of Janus-type particles further shows how anisotropic and patchy interactions can be captured analytically and extended beyond pairwise DLVO-like descriptions [GraafBoon2012, GLCBB_2025]. These results supply a compact, systematically improvable building block for modeling electrostatic interactions in colloidal and soft/biological systems, and they suggest natural extensions to more complex geometries and beyond-linear electrostatic theories. Because the expansions are ordered by screening range and interaction path length, they can be truncated with a clear physical error control, making them particularly suitable as analytical input for coarse-grained models, simulations, and the interpretation of nonadditive electrostatic effects. We anticipate that the resulting screening-ranged expansions for potentials, energies, and forces will provide a useful analytical backbone for reduced models and simulations, and for interpreting nonadditive interactions, OCR/LCA, and anisotropic colloids in experiment.
@@ I Acknowledgments
where matrix $\mathsf{K}_{ij}$ and vector $\tilde{\mathbf{S}}_{i}$ are block components of $\mathbb{K}$ and $\vec{\tilde{\mathbb{S}}}$ [supplem_pre_math]; $\tilde{\mathbf{G}}_{i}^{(0)}$ is independent of any $R_{ij}$, $\tilde{\mathbf{G}}_{i}^{(1)}$ contains elements proportional to factors $\frac{e^{-\tilde{R}_{ij}}}{R_{ij}}$ ($\tilde{R}_{ij}\mathrel{:=}\kappa R_{ij}$, $j\neq i$), $\tilde{\mathbf{G}}_{i}^{(2)}$ — to $\frac{e^{-\tilde{R}_{ik}}}{R_{ik}}\frac{e^{-\tilde{R}_{kj}}}{R_{kj}}$ ($k\neq i,j$), and so forth [supplem_pre]. Similarly, for the vector of internal coefficients $\mathbf{L}_{i}=\{L_{nm,i}\}_{0\leq\left|m\right|\leq n}$ we build
@@
where (13b) merely shows an expanded form of the $l=0,1$ addends of (13a) after explicit substitution of the Bessel functions expressions into them, $\gamma_{ji,jp}$ is the angle between $\hat{\mathbf{R}}_{ji}$ and $\hat{\mathbf{R}}_{jp}$, and $d_{0,i}\mathrel{:=}e^{2\tilde{a}_{i}}\frac{\tilde{a}_{i}-1}{\tilde{a}_{i}+1}+1$, $d_{2,i}\mathrel{:=}e^{2\tilde{a}_{i}}\frac{(\varepsilon_{i}+2\varepsilon_{\mathsf{sol}})(\tilde{a}_{i}-1)-\varepsilon_{\mathsf{sol}}\tilde{a}_{i}^{2}}{(\varepsilon_{i}+2\varepsilon_{\mathsf{sol}})(1+\tilde{a}_{i})+\varepsilon_{\mathsf{sol}}\tilde{a}_{i}^{2}}+1$. If $N=2$ and $\tilde{a}_{1}$, $\tilde{a}_{2}$ are small enough (small/weakly-screened particles) so that $d_{0,i}\approx\frac{2}{3}\tilde{a}_{i}^{3}$, $d_{2,i}\approx\frac{2(\varepsilon_{i}-\varepsilon_{\text{sol}})}{3(\varepsilon_{i}+2\varepsilon_{\text{sol}})}\tilde{a}_{i}^{3}$, then (13b) yields

### v2

@@ Abstract
We present an exact many-body framework for electrostatic interactions among $N$ arbitrarily charged spheres in an electrolyte, modeled by the linearized Poisson–Boltzmann equation. Building on a spectral analysis of nonstandard Neumann–Poincaré-type operators introduced in a companion mathematical work [supplem_pre_math], we construct convergent screening-ranged series for the potential, interaction energy, and forces, where each term is associated with a well-defined Debye–Hückel screening order and can be obtained evaluating an analytical expression rather than numerically solving an infinitely dimensional linear system. This formulation unifies and extends classical and recent approaches, providing a rigorous basis for electrostatic interactions among heterogeneously charged particles (including Janus colloids) and yielding many-body generalizations of analytical explicit-form results previously available only for two-body systems. The framework captures and clarifies complex effects such as asymmetric dielectric screening, opposite-charge repulsion, and like-charge attraction, which remain largely analytically elusive in existing treatments. Beyond its fundamental significance, the method leads to numerically efficient schemes, offering a versatile tool for modeling colloids and soft/biological matter in electrolytic solution.
@@
For a single dielectric sphere containing several point charges and located in an electrolyte (described in [Kirkwood1934]) the LPBE admits an explicit solution, but already for two interacting spheres ($N=2$) with different dielectric constants mutual polarization makes the problem highly nontrivial [Fish, Derb2, BesleyACR, LCSB, our_jcp, McClurg, Yu3]. Already in the Poisson limit of zero ionic strength, coupled multipolar expansions subject to infinitely many boundary constraints on the coefficients severely hinder analytical progress [GWJXL, our_jpcb, LCSB, BSL2014, Yu2019, DGC_softmat]. As emphasized in Ref. [BSL2014], “An implicit series expansion is known for the system of two dielectric spheres. For more than two dielectric spheres, numerical treatment is required.” Since then, image-based multiple-scattering formalisms [Qin2019, GXFQ, QPF, Freed2014], Wigner-matrix-free approaches [Yu2019, Yu3, Yu2021], and hybrid analytical-numerical schemes [GJLX, GWJXL, DGC_softmat, LSB2018, LinQS, HasStammI, Hassan_Stamm_jctc, Lindgren_jcis, FBYJBH] have substantially extended what can be done for two spheres, but a general explicit many-body ($N>2$) analytical treatment at finite ionic strength is still lacking. Ionic screening further increases the technical complexity [Fish, LLF, LinQS, DuanGan2025, Yu3], limits the applicability of image-charge constructions [DuanGan2025], and leaves key questions on like-charge attraction (LCA), opposite-charge repulsion (OCR), and asymmetric dielectric screening (ADS) largely unresolved [Yu2021, Derb2, LCSB, DGC_softmat, DuanGan2025].
@@
In this Letter, we address this gap within the LPBE/DH framework by developing an exact and explicit many-body formalism for $N$ arbitrarily charged dielectric spheres in an electrolyte. Using the spectral properties of nonstandard Neumann–Poincaré-type operators introduced in the companion mathematical work [supplem_pre_math], we construct a convergent Neumann-type expansion for the multipole coefficients of the DH potentials and thereby obtain screening-ranged series for the potentials, interaction energies, and forces. Each term in these series is associated with a definite “screening order” and contains products of DH factors $e^{-\kappa R_{ij}}/R_{ij}$ along interaction paths, providing a systematic and controllable many-body extension to the electrostatic part of the familiar Derjaguin–Landau–Verwey–Overbeek (DLVO) and to the unscreened Coulomb descriptions. The formalism applies to general fixed charge distributions, including heterogeneously charged and Janus-like particles, and yields explicit expressions that are straightforward to evaluate numerically.
@@
We illustrate the scope of our formalism by (i) deriving exact higher-order screened corrections for charged spheres, clarifying the conditions for LCA, OCR, and ADS, and (ii) constructing compact DLVO-like interaction expressions for Janus particles with controlled many-body and higher-multipole corrections. We highlight its value through two representative applications. For monopolar charged spheres, we recover the standard DLVO interaction as the leading screening order and obtain exact higher-order many-body corrections that generalize two-sphere Fisher–Levin–Li, McClurg–Zukoski, and related energy and force expressions [Fish, LLF, McClurg, FinBar2016, our_jcp, Yu3, Derb2, Derb4, BesleyACR, FilStar_jpcb]. Even in the long-studied two-sphere problem, this framework yields new analytical insight: it provides explicit conditions for the (non)occurrence of LCA and OCR, clarifies the role of dielectric contrast and ionic strength in ADS, and recovers recent image-charge-based polarization-energy results as a particular instance of the double-screened contribution [DGC_softmat, DuanGan2025, RPSA_2024, Chan]. For Janus-type particles with bipatchy surface charge, we derive compact DLVO-like expressions for effective monopole-monopole, monopole-dipole, and dipole-dipole interactions [GraafBoon2012, GLCBB_2025], and discuss how higher screening orders systematically generate many-body and higher-multipole corrections. Taken together, these examples demonstrate that the proposed screening-ranged expansions provide a versatile analytical framework for electrostatic interactions in model colloids and coarse-grained biomolecular systems, directly amenable to efficient numerical implementation. Beyond the specific cases discussed here, the framework offers a controlled many-body expansion for screened electrostatics, organized in a compact, rapidly convergent series that can underpin effective interaction models in charged soft-matter systems, from patchy colloids to biomolecular assemblies; further examples, including particles with clouds of internal point charges, higher-order fixed multipoles, or inhomogeneous free surface distributions, are developed in the joint papers [supplem_pre, supplem_pre_force].
@@
where $\vec{\tilde{\mathbb{G}}}$ collects all exterior multipole coefficients, $\vec{\tilde{\mathbb{S}}}$ gathers the source terms due to fixed charges ($\rho_{i}^{\text{f}}$, $\sigma_{i}^{\text{f}}$), $\mathbb{I}$ is the identity and $\mathbb{K}$ is a compact operator that couples different particles and multipoles through screened propagators proportional to $e^{-\kappa R_{ij}}/R_{ij}$ ($R_{ij}$ is the magnitude of vector $\mathbf{R}_{ij}=\mathbf{x}_{j}-\mathbf{x}_{i}$ connecting sphere centers). A detailed construction of $\mathbb{K}$ and $\vec{\tilde{\mathbb{S}}}$ is given in [supplem_pre_math, supplem_pre] (see End Matter). The spectral analysis in [supplem_pre_math] shows that the spectral radius of $\mathbb{K}$ considered in the relevant Hilbert space of multipole coefficients is strictly less than $1$. This guarantees that the inverse of $(\mathbb{I}+\mathbb{K})$ admits a convergent Neumann-type operator series representation,
@@
The resulting screening-ranged many-body energy and force expansions (4)-(5) converge rapidly with $\ell$ [supplem_pre, supplem_pre_force] and are easy to implement numerically. Thus, the first few addends are typically sufficient to accurately approximate the total energy $\mathcal{E}$ and force $\mathbf{F}_{i}$, and avoid the need for conventional approaches that require solving laborious systems of equations for potential coefficients or surface polarization charges. As documented in the companion works [supplem_pre, supplem_pre_force], this rapid convergence holds for representative colloidal and biomolecular parameter sets. Since the most computationally demanding step in harmonics-expansion-based approaches is solving the linear system / inverting the corresponding matrix [DY_2006], the identification of easy-to-calculate first addends of (4)-(5) already constitutes a major computational simplification. In the companion papers [supplem_pre, supplem_pre_force], we present several examples and show that including up to the double-screened addends ($\ell=2$) already captures the key features of the full energy and force profiles, including LCA and OCR.
@@
Importantly, this framework allows identifying the specific conditions for the (non)occurrence of less obvious phenomena, such as OCR and LCA. Indeed, in the particular case of two spheres less polarizable than the surrounding medium ($\varepsilon_{i}\leq\varepsilon_{\text{sol}}$) and arbitrary $q_{i}$, it turns out that odd-$\ell$ addends $\mathcal{E}^{(\ell)}\bigr|_{N=2}(R)$, such as DLVO ($\ell=1$), have the usual Coulombic behavior w.r.t. $R$, showing opposite-charge attraction and like-charge repulsion, while even-$\ell$ addends with $\ell\geq 2$ are repulsive irrespective of the charge. This structure permits OCR (through the competition between attractive odd-$\ell$ and repulsive even-$\ell$ contributions) [^2] Indeed, when $q_{1}q_{2}<0$, competition between attractive odd-$\ell$ and repulsive even-$\ell$ addends can produce short-range OCR [supplem_pre]. This corresponds to the ADS effect, an enhanced repulsion between dielectric bodies in media with high dielectric constant, observed in numerical modelling in [DY_2006, Yu2021]. but forbids LCA. However, this is not the case when particles are more polarizable than the medium ($\varepsilon_{i}>\varepsilon_{\text{sol}}$) and $\kappa>0$, where the ionic strength can play a decisive role on the occurrence of OCR and LCA. Despite the existing recent DH-level treatments on the topic (see e.g. [Yu2021] and references therein), to our knowledge, OCR in this regime has not been explicitly characterized before in the literature (see [supplem_pre] for technical details).
@@
Figure 1: a) Electrostatic potential and iso-potential contour lines obtained for a three-sphere system. $\varepsilon_{1}=2$, $\varepsilon_{2,3}=3$, $a_{1}+a_{2,3}=35$ Å, $a_{1}=10a_{2,3}$, $\varepsilon_{\text{sol}}=80$, $q_{1}=3e$, $q_{2,3}=-2e$, $\kappa^{-1}=8.07$ Å. The arrows represent the forces acting on the spheres. Spheres 22 and $3$ are identical, but located at a different distance from sphere 1. The most significant correction introduced by the inclusion of higher order screened terms consists in an excess negative potential located inside sphere 3. This potential is more pronounced in the region facing sphere 1. This induces a difference in the particle polarization, reflected by the difference in surface charge distribution shown in panel b). c) Two spheres interaction energy w.r.t. $R=R_{12}$, $R\geq 36$ Å; $a_{1}+a_{2}=35$ Å, $a_{1}=10a_{2}$, $\varepsilon_{1}=2$, $\varepsilon_{2}=3$, $q_{1}=3e$, $q_{2}=-2e$, $\varepsilon_{\text{sol}}=80$, $\kappa^{-1}=40$ Å. Lines F, 1 and 2 depict the full $\mathcal{E}^{\text{Int}}$, $\mathcal{E}^{(1)}$ (i.e. DLVO) and $\mathcal{E}^{(1)}+\mathcal{E}^{(2)}$, respectively. Lines 2.0, 2.1 and 2.1’ correspond to one-term (i.e. MZ), two-term and simplified two-term (i.e. FLL) approximations of $\mathcal{E}^{(2)}$, respectively. (Embedded inset shows a close-up view). d) Example of $\sigma_{i}^{\text{f}}$ for a generalized Janus particle: free charges $q_{i,1}$ and $q_{i,2}$ are uniformly distributed over spherical caps of polar angles $0,\,\theta_{i,1}]$ and $\pi-\theta_{i,2},\,\pi]$, respectively, while the intermediate surface (i.e. $(\theta_{i,1},\,\pi-\theta_{i,2})$) has no fixed charge.
@@
In summary, we have introduced an exact many-body framework for electrostatics of dielectric spheres in electrolytic solution within the LPBE/DH description. By exploiting the spectral properties of Neumann–Poincaré-type operators introduced in [supplem_pre_math], we obtain convergent screening-ranged expansions for potentials, energies, and forces, organized by products of DH screening factors and directly applicable to arbitrary fixed charge distributions. The formalism recovers familiar DLVO interactions at leading order and provides explicit higher-order corrections that quantify many-body and mutual-polarization effects, clarify the conditions for ADS and LCA/OCR, and connect to recent image-charge and effective-dipole results [Fish, LLF, FinBar2016, our_jcp, DGC_softmat, DuanGan2025, RPSA_2024, Chan, Qin2019, GXFQ, QPF, Freed2014]. Our treatment of Janus-type particles further shows how anisotropic and patchy interactions can be captured analytically and extended beyond pairwise DLVO-like descriptions [GraafBoon2012, GLCBB_2025]. These results supply a compact, systematically improvable building block for modeling electrostatic interactions in colloidal and soft/biological systems, and they suggest natural extensions to more complex geometries and beyond-linear electrostatic theories. Because the expansions are ordered by screening range and interaction path length, they can be truncated with a clear physical error control, making them particularly suitable as analytical input for coarse-grained models, simulations, and the interpretation of nonadditive electrostatic effects. We anticipate that the resulting screening-ranged expansions for potentials, energies, and forces will provide a useful analytical backbone for reduced models and simulations, and for interpreting nonadditive interactions, OCR/LCA, and anisotropic colloids in experiment.
@@ I Acknowledgments
where matrix $\mathsf{K}_{ij}$ and vector $\tilde{\mathbf{S}}_{i}$ are block components of $\mathbb{K}$ and $\vec{\tilde{\mathbb{S}}}$ [supplem_pre_math, supplem_pre]; $\tilde{\mathbf{G}}_{i}^{(0)}$ is independent of any $R_{ij}$, $\tilde{\mathbf{G}}_{i}^{(1)}$ contains elements proportional to factors $\frac{e^{-\tilde{R}_{ij}}}{R_{ij}}$ ($\tilde{R}_{ij}\mathrel{:=}\kappa R_{ij}$, $j\neq i$), $\tilde{\mathbf{G}}_{i}^{(2)}$ — to $\frac{e^{-\tilde{R}_{ik}}}{R_{ik}}\frac{e^{-\tilde{R}_{kj}}}{R_{kj}}$ ($k\neq i,j$), and so forth [supplem_pre]. Similarly, for the vector of internal coefficients $\mathbf{L}_{i}=\{L_{nm,i}\}_{0\leq\left|m\right|\leq n}$ we build
@@
where (13b) merely shows an expanded form of the $l=0,1$ addends of (13a) after explicit substitution of the Bessel functions into them, $\gamma_{ji,jp}$ is the angle between $\hat{\mathbf{R}}_{ji}$ and $\hat{\mathbf{R}}_{jp}$, and $d_{0,i}\mathrel{:=}e^{2\tilde{a}_{i}}\frac{\tilde{a}_{i}-1}{\tilde{a}_{i}+1}+1$, $d_{2,i}\mathrel{:=}e^{2\tilde{a}_{i}}\frac{(\varepsilon_{i}+2\varepsilon_{\mathsf{sol}})(\tilde{a}_{i}-1)-\varepsilon_{\mathsf{sol}}\tilde{a}_{i}^{2}}{(\varepsilon_{i}+2\varepsilon_{\mathsf{sol}})(1+\tilde{a}_{i})+\varepsilon_{\mathsf{sol}}\tilde{a}_{i}^{2}}+1$. If $N=2$ and $\tilde{a}_{1}$, $\tilde{a}_{2}$ are small enough (small/weakly-screened particles) so that $d_{0,i}\approx\frac{2}{3}\tilde{a}_{i}^{3}$, $d_{2,i}\approx\frac{2(\varepsilon_{i}-\varepsilon_{\text{sol}})}{3(\varepsilon_{i}+2\varepsilon_{\text{sol}})}\tilde{a}_{i}^{3}$, then (13b) (omitting $l\geq 2$-terms) yields